# PASS-FC: Progressive and Adaptive Search Scheme for Fact Checking of Comprehensive Claims

### v1

@@ 4 Experiments
FacTool-QA FELM-WK Factcheck-GPT SciFact-Open AVeriTeC X-FACT Average
Framework Base Model F1 Acc F1 Acc F1 Acc F1 Acc F1 Acc F1 Acc F1 Acc
FacTool GPT-3.5-Turbo 0.625 0.650 0.560 0.606 0.635 0.686 0.796 0.796 0.582 0.570 0.569 0.572 0.628 0.647
Factcheck-GPT GPT-4/GPT-4omini 0.755 0.818 0.665 0.763 0.710 0.674 0.588 0.560 0.576 0.562 0.316 0.246 0.602 0.604
SAFE GPT-3.5-Turbo 0.591 0.627 0.465 0.466 0.572 0.550 0.676 0.686 0.610 0.630 0.412 0.449 0.554 0.568
PASS-FC GPT-3.5-Turbo 0.672 0.785 0.586 0.704 0.698 0.757 0.728 0.733 0.666 0.658 0.618 0.623 0.661 0.710
PASS-FC GPT-4omini 0.770 0.811 0.659 0.694 0.701 0.720 0.858 0.859 0.692 0.694 0.593 0.619 0.712 0.733
Table 1: Macro-F1 and accuracy for the fact-checking task. The highest results are in bold, and the second-best results are underlined. Results within the shaded area are cited from Wang et al. (2024b), where Factcheck-GPT uses GPT-4-Turbo as the base model. We ran the remaining Factcheck-GPT results with GPT-4omini to limit costs. Additionally, we tested the GPT-4o-powered PASS-FC on Factbench (see table 4), which significantly outperforms the baselines.
@@
Datasets. To evaluate fact-checking capabilities across multiple domains, we selected six datasets for our experiments: FacTool-QA (N=233), FELM-WK (N=507), Factcheck-GPT (N=678), SciFact-Open (N=191), AVeriTec-Dev (N=500), and X-FACT (N=1,000) (Chern et al., 2023; Chen et al., 2023b; Wang et al., 2024a; Wadden et al., 2022; Schlichtkrull et al., 2023; Gupta and Srikumar, 2021). The first three datasets (FacTool-QA, FELM-WK, and Factcheck-GPT), collectively referred to as Factbench (hereafter), represent knowledge-based QA scenarios. This grouping, introduced in OpenFactCheck (Wang et al., 2024b), is based on their similar characteristics and domain focus. To be consistent with Wang et al. (2024b), We directly use the human generated atomic claims in Factbench, skipping the step of claim decomposition. SciFact-Open, AVeriTec-Dev, and X-FACT represent scientific, real-world, and multilingual fact-checking scenarios, respectively. Notably, AVeriTec and X-FACT contain the claim date as their metadata. For the other four datasets that lack this information, we used the first release date of the dataset’s paper on arXiv as their claim date. The detailed dataset descriptions is provided in Appendix A.1.
Baselines. We included the following fact-checkers as baselines: FacTool, Factcheck-GPT, and SAFE, all using evidence retrieved from Google. These models were experimented using their default settings. More details about the models and their configurations can be found in Appendix A.2.
PASS-FC Setting. To focus on developing systems deployable across multiple downstream tasks, we refrained from additional hyperparameter tuning on the test datasets. We set English as the source language for all datasets except X-FACT, which uses its source language metadata. For all experiments, unless otherwise specified, we allowed all query generation tools, retrieved top-10 evidence for each query, set the maximum iteration number to 2, and had PASS-FC reflect only on Unsupported and Inconclusive labels. We used GPT-3.5-Turbo, GPT-4-Turbo, GPT-4o, and GPT-4omini in our experiments, with temperature sets to 0.
Metrics. Consistent with previous studies ((Wang and Shu, 2023; Li et al., 2024)), we use macro-F1 and accuracy as evaluation metrics. Macro-F1 averages the F1 scores of Supported and Unsupported labels, while accuracy is calculated across all label types. This approach slightly disadvantages PASS-FC, as it can output an Inconclusive label when uncertain, a category absent in most datasets used here (See Appendix A.3 for detailed discussion). Nevertheless, PASS-FC significantly outperforms all baselines in terms of average performance (Table 1).
@@ 4.1 Main Results
We report the overall results for PASS-FC and for the baselines for few-shot fact-checking in Table 1. We have two specific observations.
PASS-FC achieves competitive results with comparable or weaker base models. The experimental results show that PASS-FC, powered by GPT-3.5-Turbo, outperforms other baseline models using the same foundation model across various datasets. When equipped with GPT-4omini, PASS-FC exhibits performance comparable to Factcheck-GPT in the first three datasets and surpasses it in overall performance across all datasets. Additionally, table 4 demonstrate PASS-FC’s strong performance on Factbench compared to baselines when equipped with GPT-4-Turbo or GPT-4o.
PASS-FC performs consistently across diverse fact-checking scenarios. The framework shows robust performance not only in the first three datasets focusing on general LLM knowledge question-answering but also in the latter three datasets representing scientific fact-checking, real-world scenarios, and multilingual fact verification. Notably, PASS-FC significantly outperforms baseline models on X-FACT and AVeriTeC datasets. This performance difference may be partially attributed to the consideration of claim timestamps, which baseline models overlook. We also show the reruned results for baselines in table 4, finding it’s challenging to reproduce the original results in the shaded area of table 1. This discrepancy likely stems from the absence of claim timestamps as metadata in the corresponding datasets, leading to verification based on the most recent information, which may have changed since the original claim was made. Further analysis of this aspect on fact-verification accuracy is in Appendix 4
4.2 Configuration for the Best Performance
We conducted a comprehensive analysis of key hyperparameters in PASS-FC. Unless otherwise specified, all experiments used GPT-4omini as the base model with default settings. Figures 4 (a) and (b) explore the impact of evidence number and reflection trigger labels, respectively. To avoid overfitting on test data and leverage the gold evidence available, we randomly selected 100 examples from the AVeriTeC training dataset for these analyses. Figure 4 (a) reveals that as the number of retrieved evidence pieces increases, the evidence recall consistently improves. However, we observed that when the evidence count exceeds 10, the F1 score begins to decline. Figure 4 (b) demonstrates that expanding the set of labels that trigger reflection consistently leads to improved performance. This finding underscores the effectiveness of the reflection mechanism in enhancing fact-checking accuracy. Figure 4 (c) illustrates the impact of iteration count across all datasets. Interestingly, we found that unlike GPT-4o, GPT-4omini does not consistently show improvement with increased iterations. This observation aligns with the general understanding that GPT-4o possesses stronger capabilities, allowing it to benefit more from reflection.
Model F1 Acc
PASS-FC 0.770 0.811
w/o Time Augmentation 0.738 0.773
w/o Advanced Search 0.752 0.785
w/o Site-restricted Search 0.749 0.783
w/o Multilingual Search 0.770 0.811
Table 2: Ablation study results for PASS-FC on the FacTool-QA dataset.
Model F1 Acc
PASS-FC 0.593 0.619
w/o Multilingual Search 0.574 0.605
w/o Language-Specific Initialization 0.571 0.599
Table 3: Ablation study results for PASS-FC on the X-FACT dataset.
@@ 4.3 Ablation Study
We conducted an ablation study to evaluate the contribution of each component in PASS-FC, using the FacTool-QA dataset with GPT-4omini as the base model under default settings in table 2. Due to the omission of the claim decomposition step for FacTool-QA, we couldn’t assess its impact. Notably, the time augmentation, which particularly includes the date metadata, significantly influenced performance. Even using the arXiv publication date (’2023-07-26’) for FacTool claims, rather than individual claim verification dates, proved effective. This suggests that a considerable portion of facts in FacTool-QA may have changed over time. Replacing advanced search with FacTool’s simpler direct questioning approach led to a performance decrease, indicating the value of advanced search even in general knowledge domains. Removing site-restriced search also led to a performance decrease, showing the importance of retrieving evidence from credible sources. The minimal impact of removing multilingual retrieval can be attributed to FacTool-QA’s focus on general domains, where English-based advanced search appears sufficient.
To further investigate the impact of PASS-FC components in a multilingual context, we conducted an additional ablation study on the X-FACT dataset using GPT-4omini in table 3. We tested two specific modifications: removing multilingual search and eliminating language-specific initialization (using English as the default source language regardless of input language). Both modifications led to a decrease in performance, with the removal of language-specific initialization having a slightly larger impact. These findings suggest that both multilingual search capabilities and language-specific initialization contribute to improved performance in multilingual fact verification tasks.
@@ 5 Conclusion
This paper presents PASS-FC, a novel fact-checking framework that uses iterative claim augmentation, advanced search, and multilingual capabilities to address challenges like temporal and entity ambiguities. Experiments show PASS-FC significantly improves fact verification across various datasets, outperforming stronger base models in general, scientific, real-world, and multilingual tasks. As online information evolves, PASS-FC will be vital for ensuring information integrity and supporting informed decisions.
@@
The framework relies solely on Google search snippets as evidence, without exploring full webpage content or following internal links. While PASS-FC is efficient, it potentially misses valuable information that could be obtained by deeper web exploration. Although such in-depth searching might significantly increase processing time, it could potentially improve fact-checking accuracy.
@@
The ablation studies are not comprehensive due to the lack of suitable benchmarks for testing individual modules. The framework’s effectiveness is primarily judged by the final fact verification results, which may not fully reflect the performance of each component. Additionally, the advanced query generation tools proposed in this work are not triggered for every claim, necessitating the testing of a large number of examples to obtain statistically significant results.
Lastly, while using Google Search offers real-time advantages, it also introduces challenges in reproducing results. The constant updates to search results make it difficult to replicate exact experimental conditions, potentially affecting the consistency and comparability of fact-checking outcomes across different time periods.
@@ A.1 Dataset Descriptions
A.2 Baseline Descriptions
@@
FactScore
FActScore (Min et al., 2023) is a baseline model for fact-checking that first uses spaCy to segment long-form text into sentences, then decomposes these sentences into atomic facts using GPT-3.5. It employs various methods to verify these atomic facts, including a no-context language model approach, a retrieve-then-verify approach using Wikipedia, and a nonparametric probability method. FActScore calculates the overall factuality of a text by aggregating the verification results of its constituent atomic facts. This structured approach to fact-checking serves as a comparison point for our PASS-FC model in evaluating the factuality of comprehensive claims.
A.3 Label Standardization

### v2

@@ 4 Experiments
Datasets. We evaluate PASS-FC on six publicly available collections: FacTool-QA, FELM-WK, Factcheck-GPT, SciFact-Open, AVeriTeC-Dev, and X-FACT (Chern et al., 2023; Chen et al., 2023b; Wang et al., 2024a; Wadden et al., 2022; Schlichtkrull et al., 2023; Gupta and Srikumar, 2021). Following Wang et al. (2024b), FacTool-QA, FELM-WK, and Factcheck-GPT are grouped as FactBench, a knowledge-oriented QA suite for which human-written atomic claims are already provided; we therefore skip claim decomposition on that subset. SciFact-Open targets scientific statements, AVeriTeC-Dev covers real-world claims, and X-FACT stresses cross-lingual verification. Detailed descriptions and statistics (Table 4) appear in Appendix A.1.
Preventing data leakage. AVeriTeC and X-FACT supply a verification date and the URLs of fact-checking sites. At query time we restrict evidence to documents published before the verification date, which should block the official fact-checking domains. A faux domains list provided in AVeriTec is also blocked. The remaining four datasets lack explicit timestamps, so we conservatively use the arXiv publication date of the dataset paper as the cut-off. FactBench and SciFact-Open are not sourced from fact-checking portals, hence pose no additional leakage risk. A discussion on search-engine reproducibility is given in Appendix A.2.
Baselines. We compare PASS-FC with four recent LLM-based fact-checkers that also rely on Google retrieval: FacTool, Factcheck-GPT, SAFE, and FIRE. Each model runs with the authors’ default hyper-parameters; descriptions (A.3.1), implementation details (Table 5) and our rationale for their selection (A.3.3) are summarized in Appendix.
Metrics. Following prior work (Wang and Shu, 2023; Li et al., 2024), we report accuracy and macro-F1, where macro-F1 averages the class-wise F1 scores of Supported and Contradicted. Because most datasets do not contain the Inconclusive label that PASS-FC can output, this metric setup mildly penalizes our system (See Appendix A.4 for detailed discussion). Aggregate results appear in Table 2.
FacTool-QA FELM-WK Factcheck-GPT SciFact-Open AVeriTeC Average
Framework Base Model F1 Acc F1 Acc F1 Acc F1 Acc F1 Acc F1 Acc
Always True - 43.2 76.0 41.5 71.0 41.0 69.6 35.6 52.9 19.6 24.4 36.2 58.8
Always False - 19.4 24.0 22.5 29.0 19.0 23.5 32.0 47.1 37.9 61.0 26.2 36.9
Random Guess - 47.3 51.5 47.3 49.9 46.6 48.1 55.0 55.0 38.2 36.8 46.9 48.3
FacTool GPT-3.5-Turbo 62.5 65.0 56.0 60.6 63.5 68.6 79.6 79.6 58.2 57.0 64.0 66.2
Factcheck-GPT GPT-4/GPT-4omini 75.5 81.8 66.5 76.3 71.0 67.4 58.8 56.0 57.6 56.2 65.9 67.5
SAFE GPT-3.5-Turbo 59.1 62.7 46.5 46.6 57.2 55.0 67.6 68.6 61.0 63.0 58.3 59.2
FIRE GPT-4o 79.0 85.1 70.0 63.9 75.5 73.7 87.4 87.4 70.7 69.2 76.5 75.9
FIRE GPT-4omini 73.0 80.9 62.0 53.0 77.0 76.1 78.9 79.1 69.6 69.2 72.1 71.7
PASS-FC GPT-3.5-Turbo 67.2 78.5 58.6 70.4 69.8 75.7 72.8 73.3 66.6 65.8 67.0 72.7
PASS-FC GPT-4o 82.8 86.3 68.2 71.2 74.0 74.6 85.7 82.2 77.6 72.0 77.7 77.3
PASS-FC GPT-4omini 77.0 81.1 66.5 70.0 70.9 72.6 85.8 85.9 69.9 70.2 74.0 76.0
- Temporal Grounding GPT-4omini 73.8 77.3 65.3 68.4 70.9 71.8 84.3 84.2 68.3 68.8 72.5 74.1
- Structured Query Generation GPT-4omini 75.2 78.5 65.3 67.8 71.2 72.3 85.3 85.3 68.7 69.6 73.1 74.7
- Credible-Source Selection GPT-4omini 74.9 78.3 66.2 69.2 70.4 72.3 82.1 82.2 69.3 69.8 72.6 74.4
- Cross-Lingual Expansion GPT-4omini 77.0 81.1 66.2 69.6 70.9 72.6 85.8 85.9 69.2 69.4 73.8 75.7
+ force XLE GPT-4omini 78.7 83.3 67.1 71.8 71.9 73.3 86.9 86.9 72.8 71.2 75.5 77.3
Table 2: Macro-F1 and accuracy for the fact-checking task. The highest results are in bold, and the second-best results are underlined. Results within the shaded area are cited from Xie et al. (2024) or Wang et al. (2024b), where Factcheck-GPT uses GPT-4-Turbo as the base model. We ran the remaining Factcheck-GPT results with GPT-4omini to limit costs. Although FIRE used Factcheck-GPT as their development set, we still retain its result here for completeness.
PASS-FC configuration. To ensure fair cross-dataset comparison, no parameter is tuned on any test split. Instead, we probe the main hyper-parameters on 100 randomly sampled AVeriTeC-train instances with GPT-4o-mini. Figure 4a varies the number of snippets kept per query: recall climbs steadily, but macro-F1 peaks at ten snippets and drops thereafter as noise accumulates. Figure 4b alters the set of verdicts that trigger reflection: performance improves whenever more labels are allowed to relaunch the search, confirming the utility of the reflection loop.
Guided by these trends we freeze one global setting for all subsequent experiments: English is taken as the source language (X-FACT uses its own language metadata); all adaptive search tools remain enabled; we keep the top-10 hits for each query; the loop is capped at two iterations; reflection fires only on Contradicted or Inconclusive. Results are reported for GPT-3.5-Turbo, GPT-4o, and GPT-4o-mini, each run with temperature = 0.01.
@@ 4.1 Main results
Table 2 compares PASS-FC with state-of-the-art few-shot fact-checkers. Token costs on AVeriTeC-Dev and SciFact-Open are presented in Table 8. Two trends are evident.
Firstly, PASS-FC is highly competitive even when it runs on weaker back-ends. With GPT-3.5-Turbo it already surpasses all baselines that use the same model. With GPT-4o-mini it matches or exceeds Factcheck-GPT and delivers higher average accuracy than FIRE (which in turn relies on a larger GPT-4o) despite FIRE being tuned on Factcheck-GPT testset. Using GPT-4o, PASS-FC attains the best mean F1 and accuracy overall.
Secondly, performance is stable across domains. The framework not only excels on the knowledge-oriented FactBench suite but also on SciFact-Open (scientific claims) and AVeriTeC (time-stamped real-world claims), where it outperforms every baseline by a wide margin—reflecting the benefit of explicit temporal grounding. Similar behaviour is observed on the multilingual X-FACT benchmark (Table 7 in the appendix): PASS-FC outperforms every baseline. Furthermore, for all frameworks, accuracy is highest for languages typologically close to English (e.g. Norwegian) and lowest for distant languages such as Serbian, mirroring the pattern reported by Huang et al. (2022b). We also compare PASS-FC with models introduced in the AVeriTeC Challenge (Schlichtkrull et al., ) in table 6 from A.5. PASS-FC is also excellent when tested under the same setting.
Figure 4c studies the effect of reflection iterations in the main experiments. GPT-4o continues to improve when allowed a second pass, whereas GPT-4o-mini gains little, confirming that stronger models can exploit the reflection loop more effectively.
Figure 5: Analysis of XLE. (a): Distribution of languages chosen by XLE in the force-XLE setting from the ablation study. (b) The distribution of reasons that choose every language.
@@ 4.2 Ablation study
The lower block of Table 2 removes each of the four novel components in turn while keeping GPT-4o-mini and all other hyper-parameters fixed.
Temporal grounding has the largest single impact (-2.0 accuracy on average); even the coarse cut-off date “2023-07-26“ used for the entire FactTool-QA set prevents many stale assertions from being judged correct. Replacing structured queries with plain natural-language questions also degrades F1, showing that advanced operators are essential for precise retrieval. Credible-source selection mainly helps on SciFact-Open, where model-selected peer-reviewed and governmental domains are crucial for reliable scientific evidence.
[…]
Testsets Query Similarity Evidence Similarity
FacTool-QA 91.2 85.6
FELM-WK 90.6 84.2
Factcheck-GPT 91.5 83.3
SciFact-Open 92.8 86.9
AVeriTeC 90.5 83.4
Table 3: Mean cosine similarity (%) between (i) XLE queries and all other queries, and (ii) XLE evidence and all other evidence, measured with the multilingual e5 encoder.
4.3 Analysis of XLE
We examine the force-XLE setting from the ablation study to understand how multilingual retrieval aids the overall pipeline.
Language distribution.
Figure 5 plots the languages chosen by XLE when it is always activated (top-4 languages are displayed). Spanish, Chinese, and French dominate, matching their global web presence.
Why each language?
For every XLE we asked GPT-4o to label the reason as Content, Source, or Diversity3.2.2). A manual audit of 100 samples confirmed $\!>\!90\%$ precision. Content accounts for the largest share, while Source and Diversity contribute roughly equally. AVeriTeC is the only benchmark that provides URL, location, and speaker metadata. Source-based triggers are common—explaining the large gain from XLE on this set.
Complementary evidence.
Does XLE merely duplicate English evidence or does it add new information? Using the official AVeriTeC gold evidence we found no multilingual snippet with cosine similarity $>0.85$ to any English gold snippet. Qualitative inspection shows that XLE often supplies parallel reports in local media (eg. the case study in Figure 7), furnishing an alternative validation path—a pattern also observed in the AVeriTeC human-study where pairwise trace overlap is below 0.25 even within the same language.
Table 3 extends the analysis to all five test sets. While XLE queries are highly similar to the monolingual ones ($\geq 90\%$), the retrieved documents are not (all $<86\%$), indicating that the search engine’s language alignment plus the heterogeneous web bring genuinely fresh evidence into the loop.
4.4 Case Study
To highlight the internal dynamics of PASS-FC, we walk through two concise examples in Appendix A.6 (Figures 6 and 7) that isolate the roles of Temporal Grounding and XLE.
In the first case, PASS-FC first attaches the year “2023,” yielding the grounded claim In 2023, the United States has 94 operating nuclear reactors.” This single token changes the outcome. In 2024 the U.S. does list 94 active reactors, [^4] https://www.eia.gov/tools/faqs/faq.php?id=207&t=21 so a system that searches immediately is likely to retrieve supportive snippets and label the claim Supportive. Grounding to 2023 steers the search toward historical statistics, from which PASS-FC retrieves evidence that only 93 reactors were operational that year, [^5] https://www.eia.gov/energyexplained/nuclear/us-nuclear-industry.php and outputs Contradicted.
The second example starts with an English-only search that returns conflicting evidence, causing the first verification attempt to stall. In the next round, XLE produces Spanish and Portuguese query variants. The Spanish query surfaces a CNN en Español article that squarely addresses the claim. This new evidence lets PASS-FC make the right verdict, illustrating how multilingual retrieval supplies otherwise unreachable support.
@@ 5 Conclusion
We introduced PASS-FC, a fact-checking framework that augments large language models with temporal grounding and an adaptive, progressively refined search scheme. Across six benchmarks—spanning general knowledge, scientific claims, real-world events, and ten languages—PASS-FC outperforms strong baselines even when running on smaller backbone models. As on-line information continues to proliferate and shift over time, methods such as PASS-FC that explicitly reason about time, provenance, and language will be essential for maintaining information integrity and supporting reliable, automated decision making.
@@ A.1 Dataset Descriptions and Statistics
A.1.1 Statistics
Detailed distribution are presented in Table 4.
Datasets Supported Contradicted Inconclusive Conflicting/Cherrypicking All
FacTool-QA 177 (76.0%) 56 (24.0%) N/A N/A 233
FELM-WK 360 (71.0%) 147 (29.0%) N/A N/A 507
Factcheck-GPT 472 (69.6%) 159 (23.5%) 47 (6.9%) N/A 631
Scifact-open 101 (52.9%) 90 (47.1%) N/A N/A 191
AVeriTeC-Dev 122 (24.4%) 305 (61.0%) 35 (7.0%) 38 (7.6%) 500
X-FACT 500 (50.0%) 500 (50.0%) N/A N/A 1000
Table 4: Datasets distributions. X-FACT is we random sampled balanced testset, with 10 languages, one containing 100 examples. For detailed description, refer to A.1
A.2 Impact of Search Engines
Google search results may change over time. This issue has been studied in fact-checking research: Chen et al. (2023a) shows that while only 30% of search output URLs overlap when queried two months apart, the veracity judgment classification remains largely stable. We’ve verified this in our experiments. Rerun our main experiment after a two-month interval on FactCheck-GPT only slightly change the result (within 0.5% on accuracy), demonstrating minimal changes from the original results. Using search engines is unavoidable for real-world fact-checking, as they represent one of the largest, continuously updated knowledge base. We consider this a limitation of the approach rather than a weakness, given the nature of real-world fact-checking tasks.
A.3 Baselines
A.3.1 Descriptions
@@
FIRE
FIRE is a fact-checking framework that iteratively integrates evidence retrieval and claim verification. It decides whether to provide a final answer or generate a new search query based on confidence in the current judgment. Notably, FIRE can make a verdict entirely based on the model’s confidence, even without external evidence.
Base Model Frameworks
gpt-4-turbo-2024-0409 PASS-FC, FactcheckGPT, FIRE
gpt-35-turbo-0613 PASS-FC, FacTool, SAFE
gpt-4omini-0718 PASS-FC, FactcheckGPT, FIRE
Table 5: Frameworks and tested base models.
A.3.2 Settings
All compared frameworks use the same version of their default base models, as detailed in the following table 5. Plus, all baseline framework settings (for instance, temperature and evidence number) used their default settings in their official codebase.
A.3.3 Rationals for Selecting Baseline Frameworks
We adopt three criteria when choosing baselines: (i) the system must target general-purpose, textual AFC; (ii) it must retrieve evidence through a public search engine rather than a closed knowledge base; (iii) it better includes a reflection module or non-trivial search heuristics, so that a fair comparison with PASS-FC is possible. All baselines we finally report satisfy these requirements.
Among the candidates in Table 1, Self-Checker (Li et al., 2024) nominally meets the rules, but its public repository offers three distinct prompt sets—one per benchmark—making it unclear which configuration reflects real-world usage; we therefore exclude it. ProgramFC (Pan et al., 2023) and KnowHalu (Zhang et al., 2024) rely on Wikipedia alone and are thus disqualified by criterion (ii). PACAR (Zhao et al., 2024) does not release code, preventing reproducibility.
The remaining candidates are HISS (Zhang and Gao, 2023), VERISCORE (Song et al., 2024), and DEFAME (Braun et al., 2024). HISS is tailored to news-domain claims and its prompts hard-code that domain, so we omit it. VERISCORE extends SAFE (Wei et al., 2024) with an upfront “trustworthy-claim” filter; because all our test sets contain only claims that require verification, this extra step is irrelevant, and we adopt the simpler SAFE. DEFAME, proposed for the AVeriTeC challenge (Schlichtkrull et al., ), is the only truly general-purpose system in that competition in our knowledge, but it is designed for mixed text-image claims and is extremely resource-intensive (around 5 h and 50\$ for 200 purely textual samples). We therefore omit it from the main comparison, although we do report a head-to-head result on AVeriTeC-DEV in the next section.
A.4 Label Standardization